OpenStudy (anonymous):

Write the standard form of the equation of a circle with endpoint of a diameter at the points (9,6) and (-5,9)

OpenStudy (anonymous):

find the radius using distance formula and dividing that answer by 2, then use x^2 + Y^2= R^2

OpenStudy (anonymous):

R being the radius!

OpenStudy (anonymous):

I have (x-5)squared + (y-15/2) squared - but I can't get the last part - what the radius is

OpenStudy (anonymous):

use distance formula.. to find diameter.. then divide by 2

OpenStudy (anonymous):

do you know what distance formula is?

OpenStudy (anonymous):

ugh, no

OpenStudy (anonymous):

Is it 10 or 100?

OpenStudy (anonymous):

i ll tell distance formula (x1 y1) and (x2 y2).. distance between them is \[\sqrt{(x1-x2)^{2}+(y1-y2)^{2}}\]

OpenStudy (anonymous):

174?

OpenStudy (anonymous):

No do \[\sqrt{(9+5)^{2}+(6-9)^{2}}\]

OpenStudy (anonymous):

\[\sqrt{205}\]?

OpenStudy (anonymous):

correct.. !

OpenStudy (anonymous):

so the answer is: (x+5)\[(x-5)^{2}+(y-15/2)^{2}=205?\]+(y-@DIV{15;2})@Sup{2}=205?

OpenStudy (anonymous):

that doesn't look right

OpenStudy (anonymous):

Noo.. what you got is the diameter.. Root(205)/2 is the radius.. so answer is ............ = 205/4

OpenStudy (anonymous):

sorry, I am more confused than ever

OpenStudy (anonymous):

how do I write the standard equation?

OpenStudy (anonymous):

(x-a)squared +(y-b)squared= (radius)squared

OpenStudy (anonymous):

you first tell me what is the co ordinate of the centre of the circle, ??? those co ordinates are (a, b).. what you got i believe is not correct

OpenStudy (anonymous):

2, 15/2